40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 21:05:23 2022, yuta, Update, BHD, AS path alignment, removing a lot of green stuff 7x
    Reply  Wed May 4 13:09:42 2022, Anchal, Update, BHD, IMC table balanced again, IMC is locking, YARM is locking 
       Reply  Wed May 4 15:08:59 2022, Tega, Update, BHD, XARM alignment to get flashing IFO_XY_ARMS_Flashing.png
          Reply  Wed May 4 18:52:43 2022, Tega, Update, BHD, POP beam too high at POP SM4 on the ITMX chamber snnp_pr2_pr3_itmy_aligment.png
             Reply  Thu May 5 14:46:22 2022, Anchal, Update, BHD, POP beam height lowered, POP_SM4 raised 
                Reply  Thu May 5 17:05:31 2022, Tega, Update, BHD, IMC & X/Y-arm alignment IFO_aligment_Y_locking_on.png
                Reply  Mon May 9 18:43:03 2022, Anchal, Update, BHD, ITMX table layout corrected 
                   Reply  Mon May 9 18:49:05 2022, Tega, Update, BHD, Relocate green TRX and TRY PDs/cams/optics from PSL table to BS table IMG_20220509_184439943.jpgIMG_20220509_184154722.jpg
                      Reply  Tue May 10 15:46:38 2022, JC, Update, BHD, Relocate green TRX and TRY components from PSL table to BS table IMG_0741.jpegIMG_0753.jpeg
                         Reply  Tue May 10 18:25:43 2022, Paco, Update, BHD, Green Transmission path 
                   Reply  Tue May 10 16:31:14 2022, Tega, Update, BHD, ITMX optlev return beam steered to QPD 
       Reply  Fri May 6 13:48:34 2022, Anchal, Update, BHD, WFS loop instability fixed 
    Reply  Thu May 12 19:55:01 2022, Tega, Update, BHD, AS path alignment AS.JPGOMCchamber.JPG
       Reply  Thu May 12 20:11:18 2022, Anchal, Update, BHD, BHDBS Output beams steered out to ITMY table signal-2022-05-12-201844.jpeg
          Reply  Thu May 12 20:24:29 2022, yuta, Update, BHD, POX and POY investigation POY.JPG
       Reply  Mon May 16 13:22:59 2022, yuta, Update, BHD, REFL and AS paths aligned at AP table REFLPOP.JPGPOPAS.JPG
          Reply  Mon May 16 19:14:17 2022, Anchal, Update, BHD, Camera set on AS path and BHDBS output path 
             Reply  Tue May 17 18:43:38 2022, Anchal, Update, BHD, Placed SRM in ITMY Chamber 
                Reply  Wed May 18 08:30:29 2022, Paco, Update, BHD, SRM OpLev 
                   Reply  Thu May 19 08:51:40 2022, Paco, Update, BHD, SRM OpLev 
                Reply  Wed May 18 09:02:52 2022, Anchal, Update, BHD, WFS1 PD centered 
                Reply  Wed May 18 17:23:15 2022, Anchal, Update, BHD, Placed PRM in BS Chamber 
                   Reply  Thu May 19 18:56:08 2022, yuta, Update, BHD, Current OSEM sensor values with all the suspensions aligned OSEMS.txt
                   Reply  Thu May 19 19:05:59 2022, Anchal, Update, BHD, BS Chamber all work finished, BHD path setup 
                      Reply  Fri May 20 20:03:48 2022, Paco, Update, BHD, ITMY chamber work finished - LO and AS overlapped Screenshot_2022-05-20_20-38-25.pngOSEMs.txt
                         Reply  Mon May 23 13:16:59 2022, Paco, Update, BHD, End of vent - checks OSEMs20220523.txtScreenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.pngPR3position.JPG
                            Reply  Tue May 24 15:21:13 2022, Anchal, Update, BHD, Freeswing tests of new SOS started 
Message ID: 16832     Entry time: Thu May 5 14:46:22 2022     In reply to: 16831     Reply to this: 16833   16837
Author: Anchal 
Type: Update 
Category: BHD 
Subject: POP beam height lowered, POP_SM4 raised 

[Anchal, JC]

We first aligned the single arm cavity resonance for both arms to get maximum flashing. As we opened the chamber, I found that the POP beam was mostly hitting the POP_SM4 mirror but was clipping about 2 mm on the top edge.

I used TT2-PR3 to lower the injection beam angle and moved pairs of ITMY-ETMY, and ITMX-ETMX to recover as much flashing as I could in the both arms. Then, I moved PR2 in pitch from 49 to 71 to maximize the arm flashing again. After these steps, the POP beam was clearly within the POP_SM4 mirror but still in the upper half of the optic and there was maybe just a mm of clearance from the top edge. I decided to raise POP_SM4 mirror by 0.14" spacer. Now the beam is still in upper half of the mirror but has a good clearance from the edge.

The POP beam is coming outside in the in-air table at as a rising beam in the nominal path near the center of the window. This beam needs to be directed to the POP camera and RFPD on the far-side of the table.

Next steps:

  • In-air table work: Setup POP camera and RFPD.
  • In ITMX chamber, rotate ITMX Oplev mirror to clear the oplev beam off POP_SM5. Change oplev beam path outside accordingly.
  • Install green transmission from X-arm steering mirrors in BS chamber.
  • Install 4 steering mirrors in ITMY chamber at the two outputs of BHD BS to direct the beam outside.
  • Figure out POX11 rotation angle and get XARM locking as well.

 

ELOG V3.1.3-